This is more of a word problem. When a number is decreased by 25% of itself, the result is 216. What is the number?
Thanks

Answer by CharlesG2(834)

This is more of a word problem. When a number is decreased by 25% of itself, the result is 216. What is the number?
(1 - 0.25)x = 216 (25% is 0.25, you are taking 100% - 25% times the number)
x - 0.25x = 216
0.75x = 216
x = 288